Many states assess the skills of their students in various grades. One program that is available for this purpose is the National Assessment of Educational Progress (NAEP). One of the tests provided by the NAEP assesses the reading skills of twelfth-grade students. In a recent year, the national mean score was 290 and the standard deviation was 38 . Assume that these scores are approximately Normally distributed, N(290 , 38 ).

Use technology to find ( ±±0.0001) the proportion of students who have scores between 346 and 383 .

.
Answer #1

Solution :

Given that ,

mean = \mu = 290

standard deviation = \sigma = 38

P(346 < x < 383) = P[(346 - 290)/ 38) < (x - \mu) /\sigma  < (383 - 290) / 38) ]

= P(1.47 < z < 2.45)

= P(z < 2.45 ) - P(z < 1.47 )

= 0.9929 - 0.9292

= 0.0637